Determine if the following ratios form a proportion. Also, write the middle terms and extreme terms where the ratios form a proportion. 
200 mL : 2.5 litre and ₹ 4 : ₹ 50

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

200 mL : 2.5 litre and₹4 :₹50
1 litre = 1000 ml
2.5 litre = 2500 ml
200 / 2500 = 2 / 5
4 / 50 = 2 / 25
Yes, these are in a proportion
Middle terms arte 2.5 litres,₹4 and Extreme terms are 200 ml,₹50

Thanks